Tuesday, July 18, 2017

Ready for run 3...

We worked hard on the presentation today and will continue that tomorrow during run 3.

Alumni visitor!

Monday, July 17, 2017

Recorded run 2

Today was the day of our second recording. We started the day by setting the Arduino to trigger at 1 cm and 8 cm, however, that resulted in almost no change from the previous recorded run. The previous had been set to 2 cm and 9 cm. This is just another problem stemming from the sensor readings, which we don't have a solution ready for.

To correct for this problem, we reset the triggers to be at 0.5 cm and 7 cm. This resulted in nearly 2.5 cm of difference at the high head in real life, and a slight change int he low head. That's not great, but at least it's different from the last experiment.

I spent the remainder of the day putting assembling the frames from the last experiment into a dramatically sped up video, set at a rate of 1 minute per second. It looks nice, but I have started recompiling the video, as the video player included with Windows is bad, and won't play it. We'll see if that's sorted out now tomorrow.

Ethan spent the day working on a Matlab program to read the water level off of each frame of the video footage we have. It's a formidable task, and will take time. Related to that task, I put more thought and work into how to detect and represent the parts of the sand that turn red from each frame of the footage we have.

I'm not sure if we'll have enough time to complete either program before we present on Friday (our last day), but we'll see.

I'm not completely certain about what Sheiny did today. She kept busy, but is currently at her job, so she is unavailable to ask.

Friday, July 14, 2017

Today we did lots of coding working on trying to find the line of the water for every picture. This has been successful for the water that had red in it, however for the clear water we are still having problems. We also started programming red dye detection in the sand.  And we measured the volumetric discharge of our pump and the volumes of our system.
This is what the computer calculated for head differences. 

Day 21: Cookout today

We had nice lunch today, and got updates from all other groups. We are now currently running tap water to flossy to hopefully flush all the red dye out. Hopefully if it all flushed out, we won't have to remove all of the sand and wash it.

We got the result from yesterday's run and we got a really nice curvy dip that signify each oscillation. We didn't spill to the floor today, which is good. 
Our successful run
Tap water to flush Flossy

Tuesday, July 11, 2017

The day Before the Storm

Today we have just been prepping for the first official test. This includes getting the buckets with pumps in them around flossy as well as mixing the solutions. We also might have gone on our last trip to Menards. We are all very excited for tomorrow.

Flossy ready for her first run.

Monday, July 10, 2017

Day 19: More Preparations

No injuries!

This morning Ethan put sand back into Flossy. It's a process that takes hours, as it's important to sprinkle the sand into water, layer by layer. This is to keep the sand more homogeneous than having dumped all of the sand in at once, and to keep air pockets out of the sand. He also started 3D printing the square dividers that will be used later. The square dividers are meant to allow for sediments of different compositions to be placed inside of Flossy.

Sheiny spent hours today cleaning the red dye out of the sand from the previous trial runs. This is very time consuming, yet she also found time to mess with the EC meter some more.

I came in late, as I had a previous commitment this morning. I took another set of data points in our calibration set up, using our new method of gathering the data. Once again, there is significant binning happening. I set up the bucket experiment I had last time to see if it made nice triangles the way it did last Friday, and those triangles were still there. It's hard to understand what's happening with this Arduino and its sensors. At this point, I think we've given up on it gather nice, clean data.

To pass the time today, the group solved 4 or 5 cases of mysterious deaths, and helped a group of gnomes figure out hat colors. All in a day's work.

I might be killed for not having labeled the axes. But this is the nice part lines from the testing in buckets.

Friday, July 7, 2017

Day 18 I think...

Today the main challenge of the day was to find out why the man ate albatross and then killed himself after. We  also cleaned out our tank and silicon-ed the middle divider to try and stop what we think is rust. We looked at the oscilloscope to try and find out what was wrong with our sensors. later we found a way for the pressure sensors to not create binned data by making it take 4 extra readings per loop and it has been working. We also did calibration test and put the data into excel to see how the sensors are working. We have also screwed Flossy's stand to a shelf  that's about three feet high to have the Arduino sit without moving. We also polyurethaned the shelf and flossy's stand.
More calibration tests

Dr. Prosise working with the oscilloscope to try and figure out whats wrong with sensors. 

Thursday, July 6, 2017

Augustana Research Get Together and Group Picture!

Image may contain: 4 people, people smiling, people standing and outdoor

Image may contain: 4 people, people smiling, table and indoor

Image may contain: food

Day 17: 2 more weeks to go

So today is a short day, we did our short presentation to Kristen Kruger (thank you!) upstairs as a trial run in the morning.

Andrew is still working on the code, trying to work out with the precision problem. Ethan is bound and determined to build a shelf for the arduino and electrical power source, since it keeps going out of place. My experiment with the metals in the salt solution continues with trial 4.

Today we are going to present with Augustana students, hopefully, it will go well, but today's lunch was lovely!

Wednesday, July 5, 2017

Day 16: Preparations

Day 16: Preparations

Today we spent our morning putting together slides for tomorrow, when groups from Augustana will be visiting. For my part, I looked at the data we gathered from Flossy last Friday, so I could have some nice plots to show. This led to the unfortunate revelation that the precision problem with our pressure transducers are worse than we had thought. Rather than giving use smooth, exponential decays in our data, we're being given saw teeth, where the transducers seem to be binning their readings in 2.5 cm chunks.

Ethan cut the PVC pipe for Flossy in an attempt to get the red dyed water to mix better. We've also added a small pump in the tank to assist in mixing the dyed water with the clear water. Ethan's slides cover the things he's built for Flossy, including the middle divide, and the gutters. Sheiny's slides were about her EC investigation that goes along with the final experiment.

To close out the day, we started another test run. We had a tense moment when we watched the water level creep toward the overflow gutter, but that tension was cut when the water pump shut off, and the water level slowly fell. On the second run though, the water level rose to a similar height, but the sand shifted enough to allow the water to quickly flush through Flossy, creating a great red mess all over the lab.


On a good note, no injuries!